solve the following expression 5(B-4)=10
solve the following equation for the variable I
5(B-4)=10
complete the following table for the equation y = x-3
Table: X | Y
-----
-1|
0|
4|

Answer by ReadingBoosters(3246)

Distribute the 5
5B - 20 = 10
5B = 10 + 20
5B = 30
B = 6
...
plug in the x values and solve
y = x-3
if x is -1, then y = -1-3 = -4
-1, -4
if x is 0, then y = 0 - 3 = -3
0, -3
if x is 4, then y = 4 - 3 = 1
4, 1
.....................
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.